Assume the total product of three workers is 120 and the total product of four workers is 160. The average product of four workers is ________, and the marginal product of the fourth worker is ________.

A) 10; 40

B) 30; 10

C) 40; 40

D) 160; 40
Answer #1

Answer

average product of n workers =total product of n worker / n
AP=TP/n
=160/4
=40 units
========
marginal product of n th worker =total output of n workers - total output of n-1 workers
MP(n)=TP(n)-TP(n-1)
MP(4)=160-120=40 units
=====
the average product and marginal product is 40 units

Option C